OpenStudy (anonymous):

For nitrogen to be a liquid, its temperature must be within 12.78 °F of –333.22 °F. The equation |x + 333.22 | = 12.78 can be used to find x, which represents the maximum and minimum temperatures at which nitrogen is a liquid. is the maximum temperature at which nitrogen is a liquid. is the minimum temperature at which nitrogen is a liquid.

OpenStudy (anonymous):

To be within 12.78 of -333.22 you have to add and subtract 12.78 from -333.22. -333.22 + 12.78 = -333.22 - 12.78 =